OpenStudy (anonymous):

The sum of two numbers is 80 and their difference is 22. Find the two numbers. show steps and check stepi had tryed and failed

OpenStudy (zarkon):

x+y=80 x-y=22 solve for x and y

OpenStudy (anonymous):

gotthat part written out

OpenStudy (paxpolaris):

using elimination method: add the two equations ....to eliminate y

OpenStudy (anonymous):

i dont know how i have so far x+y=80 x-y=22 y=x+22 x+x+22

OpenStudy (paxpolaris):

Alright using substitution method instead:\[\large x+y=80\] \[\large x-y=22\]\[\large \implies y=x+22\] now substitute \(x+22\) for \(y\) in the first equation: \[\Large \implies x+(x+22)=80\]

OpenStudy (paxpolaris):

sorry that should be \(\LARGE x-22\) \[\large y=x-22\] \[\Large \implies x+x-22=80\]

OpenStudy (anonymous):

i am still lost

OpenStudy (anonymous):

we figured that y=28 andx=52 butwe cant get the check step andhow we got it sincemy nephew can only fgureit in his head

OpenStudy (paxpolaris):

\[x-y= 22\]\[\implies x-y+y=22+y\]\[\implies x=22+y\]\[\implies x-22=22+y-22\]\[\implies x-22=y\]

OpenStudy (paxpolaris):

because \(y= x-22\), we can substitute x-22 instead of y in the first equation \(x+y=80\) becomes \(\implies \large x+x-22=80\)

OpenStudy (paxpolaris):

\[2x-22=80\]\[\implies \large2x=80+22=102\]\[\implies \Large x={102\over2}=\huge51\]

OpenStudy (paxpolaris):

plug x=51 in any of the two original equations to get y=29

OpenStudy (paxpolaris):

You can use the elimination method by adding the two equations: \[\Large\ \ x+y=80\]\[\Large\ \ x-y=22\]\[\Large \overline{ 2x+0=102}\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

the numbers are 51 and 29